Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 284     Entry time: 31 Jul 2023, 01:42
Author: Sergii, Martin 
Group: SCS 
Subject: Night shift summary 30.07.2023 

Achivements:

-We measured vertical spot size for several VFM settings

Problems:

-We cannot focus the x-ray beam with KB. The best value what we can get is ~300 um. We checked the slits of KB. The ALAS is slightly clipped on parabolic mirror of tim ing chamber. TI is not clear if this is the reason. We cannot proceed with the program for hRIXS.

-Still no signal on APD.

ELOG V3.1.4-7c3fd00